Suppose we have the following Treasury bill returns and inflation rates over an eight-year period:

Year Treasury Bills Inflation
1 9.01%         10.86%        
2 9.87            14.41           
3 7.62            8.80           
4 6.71            6.39           
5 7.20            8.56           
6 9.50            10.99           
7 12.33            15.15           
8 14.09            14.89           
a.

Calculate the average return for Treasury bills and the average annual inflation rate for this period. (Do not round intermediate calculations and enter your answers as a percent rounded to 2 decimal places, e.g., 32.16.)

b. Calculate the standard deviation of Treasury bill returns and inflation over this period. (Do not round intermediate calculations and enter your answers as a percent rounded to 2 decimal places, e.g., 32.16.)
c. What was the average real return for Treasury bills over this period? (A negative answer should be indicated by a minus sign. Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)
Answer #1

Using excel formula to calculate
a.

A B
Year Treasury Bills Inflation
1 1 9.01%     10.86%
2 2 9.87%     14.41%
3 3 7.62% 8.80%
4 4 6.71% 6.39%
5 5 7.20% 8.56%
6 6 9.50% 10.99%
7 7 12.33%            15.15%
8 8 14.09%    14.89%
Average 7.41% or 0.0741 11.26% or 0.1126
Excel formula AVERAGE(A1:A8) AVERAGE(B1:B8)
Standard Deviation 0.30% or 0.0030 3.28% or 0.0328
Excel formula STDEV(A1:A8) STDEV(B1:B8)

c. Average Real Return =(1+Average Return)/(1+Inflation)-1 =(1+7.41%)/(1+11.26%)-1 =-3.46% or -0.0346 or -0.035